### v2

# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

Abstract
This paper develops a flexible distribution-free method for collective outlier detection and enumeration, designed for situations in which the presence of outliers can be detected powerfully even though their precise identification may be challenging due to the sparsity, weakness, or elusiveness of their signals. This method builds upon recent developments in conformal inference and integrates classical ideas from other areas, including multiple testing, rank tests, and non-parametric large-sample asymptotics. The key innovation lies in developing a principled and effective approach for automatically choosing the most appropriate machine learning classifier and two-sample testing procedure for a given data set. The performance of our method is investigated through extensive empirical demonstrations, including an analysis of the LHCO high-energy particle collision data set.
Keywords: Algorithms; Conformal Inference; Machine Learning; Multiple Comparisons; Nonparametric Methods.
1 Introduction
1.1 Background and Motivation
Outlier detection is a fundamental statistical problem with numerous applications, ranging from many areas of scientific research to fraud detection and security monitoring. In the age of AI, its relevance is further growing due to concerns over the rise of potentially unrealistic synthetic data. The complexity of high-dimensional data has encouraged the use of sophisticated machine learning models for outlier detection, but these models often lack transparency and are prone to errors, complicating their reliability. This has spurred substantial interest in conformal inference (Vovk et al., 2005), which can provide principled statistical guarantees for any outlier detection algorithm under relatively mild assumptions.
While conformal inference is gaining momentum in both statistics and machine learning, its focus in the context of outlier detection has so far primarily been on individual-level outlier identification, where each data point is separately evaluated as a potential outlier. However, individual-level detection is not always feasible. In many applications, outliers may be too rare or weak to reach statistical significance (Donoho and Jin, 2015), or they may exhibit unremarkable behavior in isolation but reveal anomalous patterns when analyzed collectively (Feroze et al., 2021), such as showing under-dispersion relative to inliers.
To extend the applicability of conformal inference to these particularly challenging settings, this paper introduces a novel approach for collective outlier detection that addresses two main challenges: (1) testing the global null hypothesis that a dataset—or a subset of it—contains no outliers, and (2) estimating the number of outliers present. Our method solves these interrelated problems by leveraging in an innovative way powerful “black-box” machine learning algorithms and highly flexible, data-driven test statistics. This integrative approach is tailored to maximize power for the data at hand and offers reliable type-I error control under mild assumptions. Additionally, our method is scalable to large data sets. It builds on recent advances in conformal inference, while also revisiting and integrating several classical concepts from other areas of statistics. As we will demonstrate, this collective approach succeeds even in scenarios where individual outlier detection fails.
Our method is broadly applicable. Financial institutions, for instance, could use it to identify complex fraud schemes involving seemingly legitimate transactions. Similarly, cybersecurity systems could apply it to detect coordinated denial-of-service attacks (Ahmed and Mahmood, 2014). In high-energy physics, where collective outlier detection generally plays a central role (Vatanen et al., 2012), researchers could use our method to sift through massive datasets from particle decay sensors in the search for new particles. These signals are typically rare and weak, making individual-level detection impractical, but they present a promising use case for our approach, as previewed in Figure 1 and Section 3.2.
1.2 Contributions and Outline
We present ACODE, an Automatic Conformal Outlier Detection and Enumeration method. ACODE leverages two-sample rank tests applied to univariate conformity scores (Kuchibhotla, 2020), which can be generated by any model trained to distinguish outliers from inliers. By employing split-conformal inference, ACODE converts these scores into statistical tests controlling the type-I error rate. In addition, by utilizing the closed testing principle (Marcus et al., 1976), ACODE can construct simultaneous lower confidence bounds for the number of outliers within any subset of the test sample (Goeman and Solari, 2011). While these high-level ideas are intuitive, the flexibility of the conformal inference framework introduces substantial complexity and a great deal of implementation freedom. This brings us to the two key methodological questions addressed in this paper.
The first question involves selecting the most effective classification algorithm for computing conformity scores. We specifically examine the trade-offs between existing approaches based on one-class classifiers (Bates et al., 2023) and positive-unlabeled learning via binary classification (Marandon et al., 2024). The second question concerns the choice of rank test to determine statistical significance based on these scores. As these issues are intertwined and dependent on specific data, we present a principled, data-driven solution. The effectiveness of this approach is showcased in Figure 1, analyzing data from the 2020 Large Hadron Collider Olympics (LHCO) (Kasieczka et al., 2021).
Figure 1: Preview of performance of ACODE on the LHCO data. Left: 90% lower confidence bound for the number of outliers. Right: power against the global null hypothesis of no outliers at the 10% level (horizontal line). The results are shown as a function of the true number of outliers in a test set of cardinality 10,000. ACODE utilizes a testing procedure that may be adaptively selected (red curve) or fixed (other solid curves). Dotted curve: number of individual discoveries or power obtained by applying the Benjamini-Hochberg procedure (BH) to conformal p-values, controlling the false discovery rate below 10%.
As detailed in Section 3.2, our analysis of the LHCO data aims to (1) detect outliers, or unusual collision events, within a test sample, and (2) establish a 90% lower confidence bound for the number of these outliers. Our method not only controls errors for both tasks but also provides insightful inferences. This illustrates ACODE’s adaptability, as it selects the most suitable testing procedures in a data-driven manner. Moreover, ACODE can leverage a diverse toolbox of classification algorithms, automatically selecting an effective model for each scenario, choosing among both one-class and binary classifiers. This feature is particularly useful in practice, as the performance of different classifiers can vary substantially across scenarios, a point further elaborated in Section 3.
Additionally, Figure 1 highlights the challenge of conducting an adaptive data-driven analysis without incurring selection bias. It compares ACODE’s performance with that of a naive “cherry-picking” approach that greedily tests various classification algorithms and testing procedures, selecting the one with the most appealing outcomes. Unsurprising, such a heuristic can lead to inflated type-I errors. In contrast, ACODE is protected against selection bias while achieving similarly high power across all scenarios tested.
Figure 1 also highlights the potential advantages of an approach specifically designed for collective outlier detection over standard alternatives for individual-level discovery. The results show that ACODE leads to much more informative collective inferences compared to the Benjamini-Hochberg (BH) procedure for False Discovery Rate (FDR) control (Benjamini and Hochberg, 1995) applied to individual conformal p-values (Bates et al., 2023; Marandon et al., 2024). In fact, the BH procedure aims to identify outliers individually, and is generally not an effective solution for global testing or outlier enumeration.
This paper is structured as follows. Section 1.3 reviews relevant prior literature, Section 2 describes our method, and Section 3 demonstrates its performance on synthetic and real data. Section 4 concludes by suggesting some directions for future research. Additional methodological aspects, implementation-specific details of the method, computational shortcuts, and further numerical results are provided in the Appendices.
1.3 Related Work
Conformal inference (Vovk et al., 2005) is an active research topic, broadly seeking reliable uncertainty estimation for the predictions output by black-box machine learning models. In the context of outlier detection (Laxhammar and Falkman, 2015; Guan and Tibshirani, 2022), prior works focused on individual-level identification, often aiming to control the FDR (Bates et al., 2023; Marandon et al., 2024). By contrast, we focus on collective detection and enumeration; this tends to lead to more informative results when dealing with weak or sparse signals that are difficult to localize. This approach also connects our work to a broader literature on multiple testing and large-scale inference.
Recent advancements in large-scale inference have led to methods for signal detection and estimation of the proportion of non-null effects (Cai and Sun, 2017). Donoho and Jin (2004) proposed an elegant solution for signal detection using Tukey’s Higher Criticism statistic. Meinshausen and Rice (2006) extended the Higher Criticism method to provide a lower confidence bound for the proportion of non-null effects. Goeman et al. (2021) further advanced this approach by embedding Higher Criticism into a closed testing procedure (Marcus et al., 1976), providing simultaneous bounds for the false discovery proportion (FDP). However, the FDP bounds derived from Higher Criticism are valid only under the assumption of independent p-values, which does not hold in the conformal inference setting.
Since the works by Genovese and Wasserman (2006) and Goeman and Solari (2011), there has been significant growth in methods providing simultaneous FDP bounds (Goeman et al., 2019; Blanchard et al., 2020; Katsevich and Ramdas, 2020; Tian et al., 2023; Heller and Solari, 2023). These have been utilized across various domains, including neuroimaging (Rosenblatt et al., 2018; Goeman et al., 2023) and genomics (Ebrahimpoor et al., 2020), as well as in permutation-based (Hemerik et al., 2019; Andreella et al., 2023; Blain et al., 2022; Vesely et al., 2023) and knockoff-based (Li et al., 2024) approaches.
In this paper, we integrate conformal inference with closed testing to construct simultaneous lower confidence bounds for the number of outliers in any subset of the test set. A central challenge is selecting an effective local testing procedure suited to the data at hand. Recent findings show that a closed testing procedure is admissible if and only if all its local tests are admissible (Goeman et al., 2021), underscoring the importance of designing a closed testing approach with admissible local tests.
For local tests relying on the combination of independent p-values, Birnbaum (1954) established that there does not exist a general uniformly most powerful p-value combination method for all alternatives. Further, any (reasonable) combining method is admissible, in the sense that there exists an alternative for which it is optimal (see Heard and Rubin-Delanchy, 2018, for specific examples concerning standard combination methods).
When a uniformly most powerful test is unavailable, a possible strategy is to maximize power locally around the null hypothesis. Lehmann (1953) identified a nonparametric alternative where the Wilcoxon-Mann-Whitney (WMW) test (Wilcoxon, 1945; Mann and Whitney, 1947) is the locally most powerful rank test. Inspired by this optimality property, we will utilize the WMW rank test as a local test in the closed testing framework, although we will also consider several other options. Additionally, in Section 2.3.2, we will explore a natural extension of the WMW rank test, inspired by Shiraishi (1985).
A typical concern with closed testing is the computational cost, generally exponential. However, efficient polynomial-time shortcuts are often available (Goeman et al., 2019; Dobriban, 2020; Tian et al., 2023). In Section 2.4 and in Appendix A3 we, respectively, present the closed testing method and the computational shortcuts that make our method feasible and scalable. As we will see, different local tests often require different shortcuts.
2 Methodology
Section 2.1 starts by formally stating the problem. Section 2.2 introduces the learning component of ACODE, detailing how a classification model is trained and applied to rank the test points and a subset of the reference inliers. Section 2.3 discusses using these ranks to test local hypotheses about the presence of outliers within specific subsets of test points. Section 2.4 describes the integration of these local tests into a closed testing framework, enabling the derivation of simultaneous confidence bounds for the number of outliers in any test data subset. Section 2.5 summarizes the overall data analysis pipeline and introduces the enhancements described in Section 2.6, which make ACODE more adaptive.
2.1 Setup and Problem Statement
Consider a reference data set $Z_{1},\ldots,Z_{m}\in\mathbb{R}^{d}$ comprising $m$ observations each with dimensions $d\geq 1$, where $d$ may be large compared to $m$. We assume these are independent and identically distributed (i.i.d.) random samples from some unknown distribution $P_{0}$ on $\mathbb{R}^{d}$. Consider also a test set $Z_{m+1},\ldots,Z_{m+n}$ and, for each $j\in[n]:=\{1,\ldots,n\}$, assume $Z_{m+j}\in\mathbb{R}^{d}$ is independently sampled from some unknown distribution $P_{j}$, which may or may not be equal to $P_{0}$. That is,
[…]
In the following, we will refer to the data points sampled from $P_{0}$ as inliers and those from any $P_{j}\neq P_{0}$ as outliers. For each $j\in[n]$, define the null hypothesis $H_{j}:P_{j}=P_{0}$. Let $I_{0}:=\{j\in[n]:P_{j}=P_{0}\}$ denote the subset of inliers in the test set. Similarly, let $I_{1}=[n]\setminus I_{0}$ denote the subset of outliers in the test set.
Prior work on conformal inference for outlier detection (Bates et al., 2023; Marandon et al., 2024) has primarily focused on testing individual-level hypotheses $H_{j}$ for each $j\in[n]$, with the goal of controlling the FDR. However, as discussed in Section 1, there are many scenarios where the power to reject these individual-level hypotheses $H_{j}$ is very low, highlighting the need for an alternative approach.
In this paper, we address two key tasks related to collective outlier detection and enumeration. The first task, outlier detection, involves testing the potentially easier-to-reject composite hypothesis $H_{S}:=\cap_{j\in S}H_{j}$, which posits that a subset $S\subseteq[n]$ of test points contains no outliers. The second task, outlier enumeration, focuses on estimating the number of outliers within a subset $S\subseteq[n]$, specifically by providing a lower confidence bound, without necessarily identifying them. For both tasks, we consider cases where $S$ is either fixed or data-driven. In the latter scenario, we will utilize the closed testing principle (Marcus et al., 1976) to obtain simultaneous inferences that remain valid for all possible subsets $S$. This approach gives practitioners flexible tools for interactive data analysis.
To deal with the possibly high-dimensional nature of the data and avoid parametric assumptions about their distribution, we will tackle the aforementioned tasks using a conformal inference framework. Essentially, we will extract one-dimensional conformity scores from the data by leveraging powerful machine learning algorithms trained to distinguish outliers from inliers. Then, we will obtain distribution-free inferences by applying appropriate rank tests to these scores. With this overall strategy in view, the following section begins with a concise overview of the relevant conformal inference background.
2.2 Computing Conformity Scores and Conformal p-values
Let $(X_{1},\ldots,X_{m})$ and $(Y_{1},\ldots,Y_{n})$ indicate two vectors of conformity scores obtained by applying a function $\hat{s}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ to each calibration and test point, respectively. In general, $\hat{s}$ may depend on an independent training data set $\mathcal{D}_{\text{train}}$ and possibly also on the unordered collection of observations $\{Z_{1},\ldots,Z_{n+m}\}$. That is, for any $i\in[m]$ and $j\in[n]$,
$X_{i}=\hat{s}(Z_{i};\mathcal{D}_{\text{train}},\{Z_{1},\ldots,Z_{n+m}\}),$ $Y_{j}=\hat{s}(Z_{m+j};\mathcal{D}_{\text{train}},\{Z_{1},\ldots,Z_{n+m}\}).$ (2)
We refer to Section 2.5 for further details on the computation of the conformity scores. Here, it is important to point out that Equation (2), combined with the data generating model defined in (1), implies that the scores $(X_{1},\ldots,X_{m},(Y_{j})_{j\in I_{0}})$ are exchangeable (Marandon et al., 2024). Additionally, it is easy to see that $(X_{1},\ldots,X_{m},(Y_{j})_{j\in I_{0}})$ are mutually independent if the function $\hat{s}$ does not depend on $\{Z_{1},\ldots,Z_{n+m}\}$.
A typical approach is to leverage these scores to compute the following conformal p-value $p_{j}$ for the null hypothesis $H_{j}$, for all $j\in[n]$:
$p_{j}$ $=\frac{1}{(m+1)}\left(1+\sum_{i=1}^{m}\mathds{1}\{X_{i}\geq Y_{j}\}\right).$ (3)
As long as $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable, $p_{j}$ is a valid p-value for $H_{j}$, in the sense that $\mathbb{P}[p_{j}\leq\alpha\mid H_{j}]\leq\alpha$ for all $\alpha\in[0,1]$. Further, $p_{j}$ is not too conservative, in the sense that $\mathbb{P}[p_{j}\leq\alpha\mid H_{j}]$ is not far from $\alpha$, as long as $m$ is large and the conformity scores are almost surely distinct; the latter would be a mild assumption (since one can always add a small amount of independent noise to $\hat{s}$) but is not required.
Testing the individual hypotheses $H_{1},\ldots,H_{n}$ is a “many-to-one comparisons to a control” problem (Dunnett, 1955). Since the $n$ comparisons of $(Y_{j})$ to $(X_{1},\ldots,X_{m})$ all use the same control sample, the p-values $p_{1},\ldots,p_{n}$ are mutually dependent, complicating the application of multiple testing procedures. Bates et al. (2023) and Marandon et al. (2024) proved that $(p_{1},\ldots,p_{n})$ are positively regression dependent on a subset (PRDS) (Benjamini and Yekutieli, 2001), which implies BH controls the FDR. Further, Bates et al. (2023) and Marandon et al. (2024) have also studied the implications of such positive dependence on the validity of p-value based methods designed to test intersection null hypotheses $H_{S}$, where $S\subseteq[n]$. This is the starting point of the next section.
2.3 Local Tests for Outlier Detection
2.3.1 Existing Approaches
Consider the null hypothesis, denoted as $H_{S}$, that a fixed subset $S\subseteq[n]$ of test points contains no outliers. As outlined below, there are many tests available for this hypothesis, each with unique strengths and weaknesses. Given the limited scope of existing optimality results and the unfeasibility of predicting which method will perform best on a specific data set, ACODE is designed to integrate a toolbox of such testing procedures. It then selects the most effective approach in a principled data-driven manner (ref. Section 2.6). To decide what to include in this toolbox, we review some relevant literature below.
One approach to testing $H_{S}$ is to apply standard aggregation methods to the conformal p-values. The combination of p-values dates back to the 1930s with Fisher, Tippett, and Pearson, originally under an independence assumption (see Owen, 2009, for a review). Moving beyond independence makes the problem more challenging but is necessary in the context of conformal inference (Bates et al., 2023).
Combination methods that do not assume independence fall into three main categories: those valid for any type of dependence, which tend to be conservative in practice (Vovk and Wang, 2020); those designed for specific types of dependence, such as the PRDS assumption (Chen et al., 2023); and those based on permutations (see Stoepker et al., 2024, for a recent example involving the Higher Criticism statistic).
Combination methods with universal validity are inadmissible for conformal p-values as they are outperformed by the Simes method. Indeed, Vovk et al. (2022) demonstrated that universally valid methods are dominated by the Simes method (Simes, 1986), which is applicable to PRDS p-values (Sarkar, 2008). However, beyond scenarios featuring rare and strong signals, even the Simes method is often surpassed by alternative approaches.
As proved in Appendix A1.2, the Simes method applied to conformal p-values is inadmissible because it is dominated by the permutation-based Simes method. However, the difference between the two is minimal unless $m$ and $|S|$ are small, which makes the Simes method appealing for computational reasons.
In general, even if permutation tests are applicable, it is unclear which test statistic should be used in practice. Birnbaum (1954) demonstrated that any combination method is optimal against some alternative. For instance, Fisher’s method, recently refined by Bates et al. (2023) to accommodate the positive dependencies of conformal p-values, is well-suited for scenarios with numerous weak signals (Heard and Rubin-Delanchy, 2018).
Another elegant optimality result is given by Lehmann (1953), which shows that the WMW test is the locally most powerful rank test for the global null hypothesis $\theta=0$ under a non-parametric alternative in which the test scores $Y_{j}$ follow the mixture distribution $(1-\theta)F+\theta F^{2}$. Here, $\theta\in[0,1]$ is the proportion of outliers and $F$ denotes the cumulative distribution function of the calibration scores $X_{i}$. This model assumes the outliers have weak signals, as $F^{2}$ is the distribution of the maximum of two independent random variables with distribution $F$, where $\mathbb{P}(X_{i}<Y_{j})=2/3$ for any specification of $F$. A rare/weak nonparametric model can be obtained for very small values of $\theta$, extending the parametric model that motivates the Higher Criticism test of Donoho and Jin (2004, 2015).
As demonstrated in Appendix A1.4, the Mann-Whitney test statistic for testing $H_{S}$ can be formulated as a combination of conformal p-values: $T^{\text{MW}}_{S}=(m+1)\sum_{j\in S}(1-p_{j})$. Since $T^{\text{MW}}_{S}$ is a monotonic function of $\bar{p}_{S}=|S|^{-1}\sum_{j\in S}p_{j}$, the test statistic can be simplified to the average of conformal p-values $\bar{p}_{S}$, leading to rejection of $H_{S}$ for small values of $\bar{p}_{S}$.
The average of the p-values, also known as Edgington’s method (Edgington, 1972), has been studied by Rüschendorf (1982) and Meng (1994). They show that twice the average of the p-values is a valid p-value for arbitrary dependence. Recently, Choi and Kim (2023) demonstrated that the “twice the average” rule cannot be improved even under the assumption of exchangeability of the p-values.
In contrast, the WMW test, when utilizing the average of the conformal p-values as the test statistic, rejects $H_{S}$ if $\bar{p}_{S}\leq\alpha^{\mathrm{Perm}}$, where the critical value $\alpha^{\mathrm{Perm}}$ is determined by taking the $\alpha$ quantile of the permutation distribution of $\bar{p}_{S}$, ensuring exact validity. This critical value $\alpha^{\mathrm{Perm}}$ may be calculated either exactly in small samples or through asymptotic approximations. A well-known asymptotic approach when $m$ and $|S|$ are both large is that of Hoeffding (1948), which leads to a much more powerful test compared to the combination approach for arbitrary dependence.
In this paper, we focus on applying the ACODE method in conjunction with the Simes test, Fisher’s combination method, and the WMW test, as previewed in Figure 1. These procedures are chosen for their ease of use within our conformal inference framework and for their complementary strengths against a range of alternative hypotheses, including both sparse/strong and dense/weak effects. We detail these methods in Table LABEL:tab:standard_approaches in Appendix A1, where we discuss their underlying assumptions, precision, and the scenarios in which they are most effective. Additional, although non-exhaustive, information about existing local testing procedures can also be found in Appendix A1.
The empirical and theoretical strengths of the WMW test also motivate us to explore in the next section a flexible extension of this approach, inspired by Shiraishi (1985), that enjoys similar theoretical properties and sometimes achieves higher power in practice.
2.3.2 Shirashi’s Locally Most Powerful Rank Test
In this section, let us imagine the calibration scores $X_{1},\ldots,X_{m}$ are drawn independently from an unknown distribution $F$ and the test scores $Y_{1},\ldots,Y_{n}$ are drawn independently from a mixture distribution $(1-\theta)F+\theta G(F)$, where $\theta\in[0,1]$ is the unknown proportion of outliers and $G$ is a distribution function on $[0,1]$. As a starting point, let us assume $G$ is known. In practice, however, the test discussed in this section can be implemented using a data-driven estimate $\hat{G}$ of $G$, as detailed in Appendix A2. Importantly, we will show that this test remains valid when using $\hat{G}$ in place of $G$, provided that $\hat{G}$ is estimated in a manner that preserves the exchangeability between the calibration and test scores.
For any $S\subseteq[n]$ and any $j\in S$, let $R_{S}^{j}$ denote the rank of $Y_{j}$ among $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. For any distribution function $G$ on $[0,1]$ with density function $g$, define the statistic $T^{\text{G}}_{S}$ as:
$T^{\text{G}}_{S}$ $=\sum_{j\in S}\mathbb{E}[g(U_{m+s}^{(R_{S}^{j})})],$ (4)
where $U_{m+s}^{(R_{S}^{j})}$ is the $R_{S}^{j}$-th order statistic in a sample of Uniform random variables of size $m+s$ and $s=|S|$. The test rejects $H_{S}=\cap_{j\in S}H_{j}$ at level $\alpha\in(0,1)$ if $T^{\text{G}}_{S}$ is larger than a suitable critical value $c^{\text{G}}_{\alpha}(m,s)$, discussed later. The indicator of this rejection event is:
$\phi^{\mathrm{G}}_{S}$ $=\mathds{1}\left\{T^{\text{G}}_{S}>c^{\text{G}}_{\alpha}(m,s)\right\}.$ (5)
This reduces to the classic WMW test statistic in the special case of $G(F)=F^{2}$. Lehmann (1953) investigated the power of rank tests in the case when the distribution of the scores assigned to outliers in the test set is equal to $G(F)=F^{k}$ for any integer $k\geq 2$, which is referred to as Lehmann’s alternative. Here, $F^{k}$ represents the distribution of the maximum of $k$ independent random variables with distribution $F$. If $X\sim F$ and $Y\sim F^{k}$, then $\mathbb{P}(X<Y)=k/(k+1)$. It is worth noting that the distance of $F^{k}$ from $F$ remains the same for any specification of $F$. Lehmann (1953) proved that in the mixture model $Y_{i}\sim(1-\theta)F+\theta F^{k}$ with $k=2$, the classical WMW test that rejects for large values of
$T_{S}^{\text{WMW}}=\sum_{j\in S}R_{S}^{j}$
is the locally most powerful invariant (LMPI) test of $H_{0}:\theta=0$, i.e., it is invariant under the group of continuous and strictly increasing transformations of the scores. Shiraishi (1985) extended the result by Lehmann (1953) and proved that the locally most powerful rank test for testing $H_{0}$ given any continuous distribution function $G$ is the one in Equation (5).
Theorem 1.
[(Shiraishi, 1985)] Assume $X_{1},\ldots,X_{m}$ i.i.d. with continuous c.d.f. $F$, and $Y_{1},\ldots,Y_{n}$ i.i.d. with c.d.f. $(1-\theta)F+\theta G(F)$ for some $\theta\in[0,1]$ and some continuous c.d.f. $G$ on $[0,1]$ with density function $g$. Denote the combined sample $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$ as $(W_{1},\ldots,W_{N})$, with order statistics $W_{(1)}<\ldots<W_{(N)}$ and $N=m+n$. Denote as $R_{j}$ the rank of $Y_{j}$ in this sequence. Then, the LMPI test rejects $H_{0}:\theta=0$ for large values of:
$T^{\mathrm{G}}=\sum_{j\in[n]}\mathbb{E}[g(U_{N}^{(R_{j})})].$
Under the Lehmann’s alternative $G(F)=F^{k}$ for integers $k\geq 2$ the associated density function is $g(u)=ku^{k-1}\mathds{1}[u\in[0,1]]$ and the test statistics in Equation (4) reduce to the collection depending on $k$:
$T^{\text{WMW, k}}_{S}$ $=\frac{k}{(N+1)\cdot\ldots\cdot(N+k-1)}\sum_{j\in S}\prod_{l=0}^{k-2}(R_{S}^{j}+l),$ (6)
which are the locally most powerful rank tests against the class of Lehmann’s alternatives.
For $k=2$, the test statistic in Equation (6) differs from the WMW statistic only by a multiplicative constant. It is worth noting that the Mann-Whitney U-statistic can be expressed as $\sum_{j\in[n]}R_{j}-n(n+1)/2$, differing only by a constant from the WMW statistic. As a consequence, inference results provided by these three statistics are equivalent.
The critical value $c^{\text{G}}_{\alpha}(m,s)$ in (5) can be obtained either through a Monte Carlo simulation of the permutation distribution—assuming only the exchangeability of the null scores and feasible for small sample sizes—or via an asymptotic approximation (Shiraishi, 1985) in the limit of large sample sizes, under the additional assumption of independence.
Theorem 2.
[(Shiraishi, 1985)] In the setting of Theorem 1 the rescaled LMPI test statistic
$\frac{\sqrt{N(N-1)}(T^{\mathrm{G}}-n\mu_{N})}{\sqrt{nm\sum_{h\in[N]}(\mathbb{E}[g(U_{N}^{(h)})]-\mu_{N})^{2}}}$ (7)
with $\mu_{N}=N^{-1}\sum_{h\in[N]}\mathbb{E}[g(U_{N}^{(h)})]$ has asymptotically a standard normal distribution under the null hypothesis $H_{0}$, as $m,n\to\infty$.
As a consequence, when considering any $S\subseteq[n]$ and denoting by $s$ its cardinality, the asymptotic distribution of the test statistic $T^{G}_{S}$ is asymptotically normal with mean
$\mu_{m+s}=\frac{s}{m+s}\sum_{h\in[m+s]}\mathbb{E}[g(U^{(h)}_{m+s})]$
and variance
$\sigma^{2}_{m+s}=\frac{sm}{m+s(m+s-1)}\sum_{h\in[m+s]}\left(\mathbb{E}[g(U^{(h)}_{m+s})]-\frac{\mu_{m+s}}{s}\right)^{2}.$
We conclude this section by emphasizing that the oracle procedure studied in Theorem 2 can be translated into a practical test by replacing the unknown alternative distribution $G$ with a suitable empirical estimate $\hat{G}$, as explained in Appendix A2. As shown in Section 3, in some scenarios of interest this data-driven approximation of the oracle LMPI test can achieve much higher power compared to other local testing methods, such as Simes’, Fisher’s, and the WMW test. In any case, the true strength of our ACODE method lies in its flexibility, as it does not depend on any single testing procedure but can instead dynamically select the most effective approach based on the data at hand.
2.4 Estimating the Number of Outliers via Closed Testing
Closed testing (Marcus et al., 1976) was initially proposed within the scope of family-wise error rate control, but it is also useful to obtain simultaneous bounds for the false discovery proportion (Goeman and Solari, 2011). In this paper, we apply closed testing to construct a $(1-\alpha)$ simultaneous lower confidence bound $d(S)$ for the number of outliers in any subset $S\subseteq[n]$ of the test set. This approach is outlined by Algorithm 1. Although this procedure may seem computationally unfeasible, since it generally requires evaluating $\mathcal{O}(\exp(n))$ tests, it can be implemented efficiently in many useful cases, as discussed later.
Input : Individual hypotheses $H_{j}$, for all $j\in[n]$; a method $\phi$ for carrying out local tests $\phi_{S}$ of $H_{S}$, for any $S\subseteq[n]$; significance level $\alpha\in(0,1)$.
1
2For each $S\subseteq[n]$, test $H_{S}$ at level $\alpha$; let $\phi_{S}\in\{0,1\}$ denote the rejection indicator.
3For each $S\subseteq[n]$, adjust the local test by setting $\bar{\phi}_{S}:=\min\{\phi_{K}:K\supseteq S\}$.
4For each $S\subseteq[n]$, the lower bound for the number of true discoveries in $S$ is:
$d(S)$ $:=\min_{K\subseteq S}\{|S\setminus K|:\bar{\phi}_{K}=0\}.$ (8)
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
Algorithm 1 Closed testing blueprint for simultaneous outlier enumeration
The first step of Algorithm 1 tests all possible intersection hypotheses $H_{S}$. We refer to these tests as the local tests. In the second step, a multiplicity adjustment is applied to all local tests. The latter results in $H_{S}$ being rejected, setting $\bar{\phi}_{S}=1$, if and only if $H_{K}$ was locally rejected for all supersets $K$ of $S$. Finally, $d(S)$ is given by (8). This is easiest to explain in the special case of $S=[n]$, when the lower confidence bound for the total number of outliers is given by the difference between the total number of test points, $n$, and the cardinality of the largest subset for which the multiplicity-adjusted local test failed to reject the null hypothesis. For example, if the global null $H_{[n]}$ is rejected but there exists a subset $K\subseteq[n]$ with cardinality $|K|=n-1$ for which $\phi_{K}=0$, then $d([n])=1$.
The following result, due to Goeman and Solari (2011), guarantees that closed testing leads to simultaneously valid confidence bounds as long as the local tests are valid.
Proposition 1 (Goeman and Solari (2011)).
If Algorithm 1 is applied using a valid local test $\phi$ satisfying $\mathbb{P}[\phi_{I_{0}}=1\mid H_{I_{0}}]\leq\alpha$, then the output lower bounds $d(S)$ satisfy:
$\mathbb{P}\left[d(S)\leq|I_{1}\cap S|\text{ for all }S\subseteq[n]\right]\geq 1-\alpha.$ (9)
The availability of computational shortcuts for Algorithm 1 hinges on the choice of local tests (Goeman et al., 2019, 2021; Tian et al., 2023). See Appendix A3 for a review of relevant shortcuts and implementation details. Appendix A3.1 reviews the shortcut of Goeman et al. (2019), which reduces the cost of Algorithm 1 to $\mathcal{O}(n\log n)$ in the case of local Simes tests. Appendix A3.3 presents a shortcut for an adaptive version of the Simes local test. Appendix A3.4 introduces a shortcut for Shirashi’s LMPI test described in Section 2.3.2. Appendix A3.5 reviews the strategy of Tian et al. (2023), which provides similar shortcuts for a family of local tests including Fisher’s method and the WMW test.
2.5 Conformalized Closed Testing
We now present Algorithm 2, which summarizes the analysis pipeline described above, starting from the training of a classifier. In short, this applies the closed testing method defined in Algorithm 1 based on local tests that take as input the calibration and test scores, assuming a practical computational shortcut is available. Therefore, Algorithm 2 outputs a simultaneous lower bound $d(S)$ guaranteed to satisfy (9).
Valid conformity scores can be obtained through various methods. A common approach involves using a one-class classifier trained on an independent set of inliers (Bates et al., 2023). Alternatively, one can employ a binary classifier via positive-unlabeled learning (Marandon et al., 2024). Another option arises when labeled outliers are available; here, the scores can be generated using a binary classifier as discussed in Liang et al. (2024). For simplicity, we focus on scenarios where all labeled data consist of inliers.
Input : Inlier data $\mathcal{D}^{\text{train}}=\{Z^{\text{train}}_{1},\ldots,Z^{\text{train}}_{m_{\text{train}}}\}$ and $D^{\text{cal}}=\{Z_{1},\ldots,Z_{m}\}$.
Test data $D^{\text{test}}=\{Z_{m+1},\ldots,Z_{m+n}\}$. Significance level $\alpha\in(0,1)$.
Machine learning algorithm $\mathcal{A}$ for one-class or binary classification.
Chosen local testing method $\phi$; e.g., Simes, WMW, etc.
1
2if $\mathcal{A}$ is a one-class classification algorithm then
3       Train $\mathcal{A}$ using the data in $(Z^{\text{train}}_{1},\ldots,Z^{\text{train}}_{m_{\text{train}}})$.
4
5 else if $\mathcal{A}$ is a binary classification algorithm then
6       Train $\mathcal{A}$ using $(Z^{\text{train}}_{1},\ldots,Z^{\text{train}}_{m_{\text{train}}})$ and $\{Z_{1},\ldots,Z_{m},Z_{m+1},\ldots,Z_{m+n}\}$.
7
8 Apply $\mathcal{A}$ to evaluate the calibration and test scores, $(X_{1},\ldots,X_{m})$ and $(Y_{1},\ldots,Y_{n})$.
9
10For any desired $S\subseteq[n]$, compute $d(S)$ using Algorithm 1 (ref. Appendix A3).
11
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
Algorithm 2 Conformal Outlier Detection and Enumeration
Proposition 1 implies Algorithm 2 produces valid inferences as long as all local tests are valid, and our method is precisely designed to achieve this (ref. Section 2.3). While the lower bound $d(S)$ output by Algorithm 2 is simultaneously valid for all possible subsets $S\subseteq[n]$, in the sense of (9), this does not mean Algorithm 2 needs to explicitly output $d(S)$ for all $S\subseteq[n]$. On the contrary, one would typically apply Algorithm 2 focusing on a particular (but possibly data-driven) choice of $S$, as demonstrated in Section 3. In any case, the components of Algorithm 2 involving the model training and the computation of the conformity scores only need to be applied once, irrespective of $S$.
We conclude this section by emphasizing that the flexibility of Algorithm 2, which can accommodate a variety of classifiers and testing procedures, introduces some challenges. In particular, different classification algorithms and testing procedure may result in significant performance variation across different data sets, and it is unclear a priori how to maximize power. Unfortunately, leaving too much latitude to practitioners is not always desirable, as it may inadvertently encourage “cherry picking” behaviors that, as previously illustrated in Figure 1, can result in invalid inferences. This issue motivates the extension introduced in the next section. This extension will channel the flexibility of Algorithm 2 into a principled, automatic method that often achieves competitively high power in practice while adding a layer of protection against the risks of human-driven selection bias.
2.6 Data-Driven Tuning
We extend Algorithm 2 to leverage a potentially diverse suite of classifiers and local testing procedures. The goal is to approximately maximize power while controlling the type-I errors. Our solution, outlined by Algorithm 3, is inspired by similar approaches proposed by Liang et al. (2024) and Marandon et al. (2024) in the context of individual outlier identification under FDR control. The idea is to embed the data-driven tuning step within an additional layer of sample splitting. As detailed below, this is crucial for maintaining the validity of local tests after selecting the most effective classifier and local testing method.
Input : Inlier data $\mathcal{D}^{\text{train}}=\{Z^{\text{train}}_{1},\ldots,Z^{\text{train}}_{m_{\text{train}}}\}$, $D^{\text{cal}}=\{Z_{1},\ldots,Z_{m}\}$, $\mathcal{D}^{\text{tune}}=\{Z^{\text{tune}}_{1},\ldots,Z^{\text{tune}}_{m_{\text{tune}}}\}$.
Test data $D^{\text{test}}=\{Z_{m+1},\ldots,Z_{m+n}\}$. Significance level $\alpha\in(0,1)$.
A list of algorithms $\mathcal{A}^{(1)},\ldots,\mathcal{A}^{(K)}$ for one-class or binary classification.
A list of local testing methods $\phi^{(1)},\ldots,\phi^{(L)}$; e.g., Simes, WMW, etc.
1
2Define $\tilde{\mathcal{D}}^{\text{test}}=\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$.
3 for each $k\in[K]$ do
4
5      for each $l\in[L]$ do
6             Compute $d_{k,l}(S)$ by applying Algorithm 2 based on $\mathcal{A}^{(k)}$ and $\phi^{(l)}$, using $\mathcal{D}^{\text{train}}$ for training, $\mathcal{D}^{\text{tune}}$ for calibration, and $\tilde{\mathcal{D}}^{\text{test}}$ as the test set.
7             Compute $\hat{p}_{k,l}(S)$, the p-value for $H_{S}$ based on $\mathcal{A}^{(k)}$ and $\phi^{(l)}$.
8
9
10Find $(\hat{k}(S),\hat{l}(S))=\arg\min_{k\in[K],l\in[L]}\arg\max_{k\in[K],l\in[L]}d_{k,l}(S)$; if not unique, look at $\hat{p}_{k,l}(S)$.
11
12Compute $d(S)$ by applying Algorithm 2 based on $\mathcal{A}^{(\hat{k}(S))}$ and $\phi^{(\hat{l}(S))}$, using $\mathcal{D}^{\text{train}}\cup\mathcal{D}^{\text{tune}}$ for training, $\mathcal{D}^{\text{cal}}$ for calibration, and $\mathcal{D}^{\text{test}}$ as the test set.
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
13
Algorithm 3 Automatic Conformal Outlier Detection and Enumeration
The initial tuning module of Algorithm 3 sees the calibration and test data only through the lenses of the unordered collection $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$. Consequently, when Algorithm 2 is applied again in the inference step of Algorithm 3, using $(Z_{1},\ldots,Z_{m})$ as a calibration set and $(Z_{m+1},\ldots,Z_{m+n})$ as a test set, the scores $(Z_{1},\ldots,Z_{m+n})$ corresponding to inliers are still exchangeable conditional on the selected classifier and testing procedure. Thus, Algorithm 3 can often achieve high power without selection bias, as previewed in Figure 1 and confirmed in Section 3. Formally, Algorithm 3 also outputs a simultaneous lower bound $d(S)$ guaranteed to satisfy (9), as long as the local tests are valid.
There is an interesting distinction between exchangeability and independence in the context of Algorithm 3. Recall from Section 2.2 that the model in (1) and the approach described in (2) lead to i.i.d. scores for the inliers in the case of one-class classification, and only exchangeable scores in the case of binary classification (Marandon et al., 2024).
Algorithm 3 changes this picture a little. As we condition on the selected classifier and closed testing procedure, the scores of the inliers among $(Z_{1},\ldots,Z_{m+n})$, although still exchangeable, can no longer be independent, irrespective of whether they were initially obtained through one-class or binary classification. This is because such conditioning is similar to (but weaker than) conditioning on the unordered collection $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$, an operation that transforms the distribution of $(Z_{1},\ldots,Z_{m+n})$ into a finite-population distribution. However, this issue does not seem to have noticeable practical implications. This robustness can be explained by noting that, among the local testing procedures considered in this paper, independence is only needed to derive the large-sample asymptotic distribution of the WMW and Shirashi test statistics. (In small samples, their critical values can be obtained via permutations, under the sole assumption of exchangeability.) In large samples, however, conditioning on $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$ makes little practical difference because the cardinality of this set is large, since all our scores always take distinct values in practice.
3 Empirical Demonstrations
Section 3.1 applies ACODE to synthetic data, and Section 3.2 to the 2020 LHCO data (Kasieczka et al., 2021). Additional experiments are described in Appendix A4.
3.1 Experiments with Synthetic Data
3.1.1 Setup
We apply ACODE on synthetic data from a distribution inspired by Liang et al. (2024) and Bates et al. (2023). Each observation $Z_{i}\in\mathbb{R}^{1000}$ is sampled from a multivariate Gaussian mixture $P_{Z}^{a}$, such that $Z_{i}=\sqrt{a}\,V_{i}+W_{i}$, for some constant $a\geq 0$ and vectors $V_{i},W_{i}\in\mathbb{R}^{1000}$. The inliers correspond to $a=1$ and the outliers to $a=0.7$. The elements of $V_{i}$ are standard Gaussian, while each element of $W_{i}$ is independent and uniformly distributed on a discrete set $\mathcal{W}\subseteq\mathbb{R}^{1000}$ with $|\mathcal{W}|=1000$. The vectors in $\mathcal{W}$ are independently sampled from Uniform($[-3,3]^{1000}$) prior to the first experiment. The numbers of inliers in the training, calibration, and tuning sets utilized by ACODE are 1000, 750, and 250, respectively.
ACODE is applied using 6 classification algorithms and 5 local testing procedures. The algorithms include 3 one-class classifiers (isolation forest, support vector machine, and “local outlier factor” nearest neighbors) and 3 binary classifiers (deep neural network, random forest, and AdaBoost), all implemented in the Python package scikit-learn. The one-class classifiers compute out-of-sample scores, e.g., as in Bates et al. (2023), while the binary classifiers compute in-sample scores with the approach of Marandon et al. (2024).
The testing procedures leveraged by ACODE are: Simes’ test with and without Storey’s correction, Fisher’s combination method, the WMW test, and the LMPI test described in Section 2.3.2, applied using $G(F)=F^{3}$. Recall that this choice of $G$ is optimal against Lehmann’s alternative with $k=3$, which may or may not fit our data well. An implementation of the LMPI test based on a data-driven estimate of $G$ will be considered later.
3.1.2 Global Outlier Enumeration
We begin by constructing 90% lower confidence bounds for the total number of outliers in a test set of size 1000, varying the proportion of outliers as a control parameter. Figure 2 summarizes median lower bounds produced by ACODE over 100 independent experiments, separately for different classifiers and local testing procedures. The left and center panels compare the performance of ACODE applied using only the 3 one-class or binary classifiers, respectively, while the right panel corresponds to ACODE automatically selecting a classifier from the full suite of all 6 options.
Figure 2: Median values for a 90% lower confidence bound on the number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. The results are shown as a function of the true number of outliers within a test set of size 1000. The most adaptive version of ACODE can automatically select an effective classifier and local testing procedure in a data-driven way.
The findings highlight that one-class classifiers generally outperform binary classifiers on these data, consistently with the observations of Liang et al. (2024), although binary classifiers can be advantageous in other contexts, as shown in Appendix A4.1. Figure 2 also distinguishes between the distinct performances of the 5 local testing procedures (shown in different colors), and the scenario in which ACODE selects one procedure adaptively. For these data, the WMW test yields the most informative (highest) lower confidence bounds, and ACODE’s performance closely approximates that of an ideal oracle that knows the optimal combination of classifier and testing procedure.
3.1.3 Selective Outlier Enumeration
Figure 3 reports on related experiments in which the goal is to construct a 90% lower confidence bounds for the number of outliers within a data-driven subset of test points, selected as those with the largest conformity scores. To facilitate the interpretation of these experiments, ACODE is applied using a fixed classification algorithm (a one-class support vector machine), not a suite of 6 different algorithms. This ensures the test subsets are always selected based on comparable conformity scores in all repetitions of the experiments. These results further demonstrate how the efficacy of different local testing procedures can vary in different situations. Simes’ method performs better with very small selected sets, while the LMPI test with $G(F)=F^{3}$ excels in cases with moderately large selected sets. Once more, ACODE can approximately maximize power by autonomously identifying the most effective classifier and local testing procedure for each scenario.
Figure 3: Median values for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in experiments similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve corresponds to the true number of outliers in this selected set. In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
3.1.4 Hunting for Adversarial Anomalies
We now apply ACODE leveraging an additional local testing procedure, utilizing a more flexible implementation of the LMPI approach outlined in Section 2.3.2. This approach enables our method to even detect elusive anomalies hidden by an adversary.
We consider inlier data sampled from a standard multivariate normal distribution with 100 independent components. The outliers are generated by an “adversary” agent through the following process. The adversary first trains a one-class support vector machine on a separate dataset consisting of 1,000 inlier points. Then, to generate each outlier, it randomly produces three independent inliers and selects the one closest to the decision boundary of the support vector machine. This approach results in outliers that are difficult to detect on an individual basis but can be identified collectively, as their conformity scores tend to be under-dispersed compared to those of the inliers.
Without prior knowledge of the distribution of outlier scores, ACODE is applied as in previous experiments, using the same toolbox of six machine learning classifiers, but including a sixth local testing procedure. This procedure is a practical approximation of the LMPI test described in Section 2.3.2, using an empirical estimate $\hat{G}$ of $G$. This estimate is obtained by fitting a mixture of Beta distributions, while preserving the exchangeability between calibration and test scores, as detailed in Appendix A2.
The results reported in Figure 4 demonstrate that our method effectively detects these elusive adversarial outliers, even though the first five local testing procedures considered become ineffective in this scenario. For additional experiments that offer a more detailed view of the performance of our empirical approximation of the LMPI local testing procedure in different settings, please refer to Figures A6A7 in Appendix A4.1.
Figure 4: Performance of ACODE on synthetic data with adverserially hidden outliers exhibiting underdispersed conformity scores. Most local testing procedures cannot detect these outliers, but a data-driven approximation of the LMPI test enables our method to achieve high power. Other details are as in Figure 2.
3.2 Experiments with High-Energy Particle Collision Data
We now apply our method on data from the 2020 LHCO contest (Kasieczka et al., 2021), revisiting the example previewed in Figure 1. Participants in this contest were asked to estimate the number of interesting “signal” particle collision events within a given data set by contrasting it with a “reference” data exclusively comprised of “background” events generated by a physics simulation software. While the LHCO challenge is closely related to the problem considered in this paper, the analysis in this section overlooks some of the physics intricacies and emphasizes instead the statistical aspects of our method.
We focus on a subset of the LHCO data containing 1,100,000 observations of 14 high-level features constructed based on subject domain knowledge. Each observation was labeled as either a background (inlier) or a 2-prong signal (outlier) event, where the latter refers to hadron collision resulting in the emission of two charged particles. The goals are to detect and enumerate the outliers within a randomly chosen unlabeled test set, having access to a reference set of independent and identically distributed inliers.
In each experiment, we randomly select a test set of 10,000 events, among which the proportion of outliers is a control parameter varied between 0 and 0.15. We also randomly sample three disjoint subsets of inliers, to be used for training (cardinality 10,000), calibration (cardinality 1000), and tuning (cardinality 1000), respectively. ACODE is then applied as described in Section 3.1, leveraging the same suites of 6 classification algorithms and 5 local testing procedures. The additional local testing procedure described in Section 3.1.4 (the data-driven approximation of the LMPI oracle) is omitted here, for simplicity. This omission is due to its somewhat higher computational cost of its closed testing shortcut and the lack of clear advantages observed with these data.
Figure 1 compares the performance of different implementations of our method, in terms of power and the total number of detected outliers, as a function of the true number of outliers. We refer to Figure A8 and Table A3 in Appendix A4.1 for a more comprehensive view of these results. The latter detail the performance of ACODE separately applied based on each of the 5 alternative local testing procedures considered here— to prevent overcrowding, only a subset of these local procedures were previously displayed in Figure 1.
Note that these results also include comparisons of ACODE’s performance to that of a naive “cherry-picking” heuristic version of our method, which does not provide valid inferences, as well as to that of the BH procedure applied to conformal p-values. The latter is an effective approach for individual outlier identification under FDR control (Bates et al., 2023; Marandon et al., 2024), but it is not designed to estimate the number of outliers.
Overall, ACODE yields more informative inferences compared to individual-level outlier identification, especially when leveraging the WMW test or Fisher’s combination method. Further, ACODE achieves near-oracle performance with respect to the selections of classifier and local testing procedure, without incurring in selection bias.
Figure 5 describes related results from experiments in which ACODE is applied to construct a 90% lower confidence bound for the number of outliers in a data-driven subset of test points, selected as those with the largest conformity scores. Similar to Figure 3, in these experiments the scores are computed by a single classifier, AdaBoost, to facilitates the interpretation of the findings. The results confirm the LMPI test for Lehmann’s alternative with $k=3$ tends to lead to more informative lower bounds compared to other local testing procedures, including the WMW test. Moreover, the adaptive version of ACODE again approximately maximizes power by selecting the testing procedure automatically. Additional results, with qualitatively similar conclusions, are presented in Appendix A4.2.
Figure 5: Median values for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve corresponds to the true number of outliers in this selected set.
4 Discussion
This work opens several avenues for future research. For example, while our method employs sample splitting, one could explore extensions to more sophisticated conformal inference frameworks that use cross-validation, which tends to be more computationally expensive but also more powerful for small sample sizes (Barber et al., 2021). Additionally, this paper assumes an ideal reference data set of inliers; future work could examine relaxations of this assumption, accounting for possible contamination of the calibration data (Barber et al., 2023), or other distribution shifts (Tibshirani et al., 2019) following a weighting approach similar to that of Hu and Lei (2023).
Lastly, an intriguing question is whether the optimality of local tests extends to closed testing procedures, a concept intertwined with the notion of admissibility. Although this issue remains unexplored, a starting point may be found in Goeman et al. (2021), which provides a general discussion on admissibility and its implications for closed testing.
Software Availability
An open-source software implementation of the methods described in this paper, and code needed to reproduce all results, are available online at https://github.com/msesia/conformal_nout.
Appendix A1 Background on Local Testing
Method Validity Precision Power References
Simes PRDS Exact if $\alpha(m+1)/|S|$ is an integer Rare/Strong effects Simes (1986), Sarkar (2008)
Fisher Asymptotic§ (Permutations) Conservative (Exact) Dense/Weak effects Fisher (1925), Bates et al. (2023)
WMW Permutations Exact Rare/Weak effects Wilcoxon (1945), Mann and Whitney (1947)
A1.1 The Simes Test for Conformal p-Values
The Simes combining function is a test statistic based the ordered $p$-values:
$T^{\mathrm{Simes}}_{S}$ $=\min_{k\in\{1,\ldots,|S|\}}|S|\cdot\frac{p_{(k:S)}}{k},$ (A10)
where $p_{(k:S)}$ denotes the $k$-th smallest value in the multiset $\{p_{j}:j\in S\}$. The Simes combining function is symmetric because it remains unchanged under any permutation of $(p_{j})_{j\in S}$.
The Simes test (Simes, 1986) rejects $H_{S}$ at level $\alpha\in(0,1)$ if and only if $\phi^{\mathrm{Simes}}_{S}=1$, where
$\phi^{\mathrm{Simes}}_{S}$ $=\mathds{1}\left\{T^{\mathrm{Simes}}_{S}\leq\alpha\right\}.$ (A11)
The PRDS property of the conformal $p$-values in (3) implies that the Simes test applied to the conformal $p$-values $(p_{j})_{j\in S}$ is a valid level-$\alpha$ test for the intersection null hypothesis $H_{S}$ (Sarkar, 2008): for any given $S\subseteq[n]$, $\mathbb{P}(\phi^{\mathrm{Simes}}_{S}=1|H_{S})\leq\alpha$.
We say that a level-$\alpha$ test $\phi_{S}$ dominates another level-$\alpha$ test $\psi_{S}$ if $\phi_{S}\geq\psi_{S}$. The domination is strict if, in addition, $\phi_{S}>\psi_{S}$ for at least one $(p_{j})_{j\in S}$. We say that a test is admissible if it is not strictly dominated by any test.
Theorem 3.1 in Vovk et al. (2022) demonstrates that the Simes test dominates all tests valid under arbitrary dependence of the p-values and relying on symmetric combining functions.
A1.2 The Permutation-Based Simes Test
The Simes test for conformal $p$-values is itself inadmissible, as it is dominated by the permutation-based Simes method.
The theory of permutation tests and related references can be found in the books Cox and Hinkley (1979), Lehmann and Romano (2005), Pesarin and Salmaso (2010), as well as in the works of Hemerik and Goeman (2018, 2021).
Write $(W_{1},\ldots,W_{m+|S|})$ for the combined sample $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. Consider the permutation group $\Pi$ whose elements are permutations of the set of integers $[m+|S|]$. Choose a test statistic $T=T(W_{1},\ldots,W_{m+|S|})$ for which small values are to be regarded as evidence against $H_{S}$. For any $\pi\in\Pi$, let $T_{\pi}$ denote the value of $T$ applied to the permuted vector $(W_{\pi(1)},\ldots,W_{\pi(m+|S|)}))$. For $\alpha\in(0,1)$, the permutation critical value is defined by:
$\alpha^{\mathrm{Perm}}=\max\left(a\in\{0,(T_{\pi})_{\pi\in\Pi}\}:\sum_{\pi\in\Pi}\mathds{1}\{T_{\pi}\leq a\}\leq\alpha|\Pi|\right).$ (A12)
Then, the permutation-based Simes test with $T=T_{S}^{\mathrm{Simes}}$ in (A10) rejects $H_{S}$ at level $\alpha$ if and only if $\phi^{\mathrm{SimesPerm}}_{S}=1$, where
$\phi^{\mathrm{SimesPerm}}_{S}$ $=\mathds{1}\left\{T^{\mathrm{Simes}}_{S}\leq\alpha^{\mathrm{Perm}}\right\}.$ (A13)
A permutation test is exactly valid, i.e. $\mathbb{P}[\phi^{\mathrm{Perm}}_{S}=1\mid H_{S}]=\alpha$, for any choice of test statistic under the sole assumption that the scores $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable. However, achieving exact validity for all values of $\alpha$ is generally not feasible due to the discreteness of the permutation distribution. Nevertheless, it is possible to ensure exactness through a mathematical artifice using a randomized critical region.
The computational cost of a permutation test may be $\mathcal{O}((m+|S|)!)$, which is prohibitive if $m$ or $|S|$ are even moderately large. The typical solutions involve restricting the set of permutations, either by utilizing a fixed subgroup or a random subset of $\Pi$ (chosen independently of the conformity scores) with an added trivial identity permutation; e.g., see Theorems 1 and 2 in Hemerik and Goeman (2018).
The precise computation of $\alpha^{\mathrm{Perm}}$ for the Simes statistic requires (${n+|S|\choose|S|}$ permutations, representing the number of ways, disregarding order, that $|S|$ test units can be chosen from among $n+|S|$ units. However, this $\alpha^{\mathrm{Perm}}$ does not depend on the scores $(W_{1},\ldots,W_{m+|S|})$ but only on $m$, $|S|$ and $\alpha$, making it possible to tabulate. This is because the permutation-based Simes test is a rank test.
To see this, consider the conformal $p$-value corresponding to the permuted vector $(W_{\pi(1)},\ldots,W_{\pi(m+|S|)}))$, that is:
$p^{\pi}_{j}$ $=$ $\frac{1}{(m+1)}\left(1+\sum_{k=1}^{m}\mathds{1}\{W_{\pi(k)}\geq W_{\pi(m+j)}\}\right).$
This yields $p_{j}$ in (3) for the identity permutation. Note that $p^{\pi}_{j}$ does not depend on the scores $(W_{1},\ldots,W_{m+|S|})$, but only on their ranks.
The next proposition demonstrates that the Simes test based on conformal p-values is inadmissible, as it is strictly dominated by the permutation-based Simes test. Moreover, the two tests coincide when $\alpha(m+1)/|S|$ is an integer.
Proposition A1.
The $\alpha$-level test $\phi^{\mathrm{SimesPerm}}_{S}$ in (A13) dominates the $\alpha$-level test $\phi^{\mathrm{Simes}}_{S}$ in (A11). The domination is strict for at least one combination of $\alpha$, $m$, $n$ and $S\subseteq[n]$. Moreover, it holds that
$\frac{|S|}{m+1}\left\lfloor\alpha\frac{m+1}{|S|}\right\rfloor\leq\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\mid H_{S}]\leq\mathbb{P}[\phi^{\mathrm{SimesPerm}}_{S}=1\mid H_{S}]\leq\alpha,$ (A14)
and if $\alpha(m+1)/|S|$ is an integer, then all inequalities in (A14) becomes equalities, i.e. the Simes test based on conformal p-values is of exact size $\alpha$.
Proof.
If $\alpha^{\mathrm{Perm}}>\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}\geq\phi^{\mathrm{Simes}}_{S}$.
If $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}\leq\phi^{\mathrm{Simes}}_{S}$. We need to show that if $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}=\phi^{\mathrm{Perm}}_{S}$ for all $(p_{j})_{j\in S}$. We will derive a contradiction. Suppose $\phi^{\mathrm{Simes}}_{S}>\phi^{\mathrm{PermSimes}}_{S}$ for some $(\tilde{p}_{j})_{j\in S}$. The corresponding test statistic $\tilde{T}^{\mathrm{Simes}}_{S}$ must be $\alpha^{\mathrm{Perm}}<\tilde{T}^{\mathrm{Simes}}_{S}\leq\alpha$. Since $\tilde{T}^{\mathrm{Simes}}_{S}=T_{\tilde{\pi}}$ for some $\tilde{\pi}\in\Pi$, it follows that $|\Pi|^{-1}\sum_{\pi\in\Pi}\mathds{1}\{T_{\pi}\leq\alpha\}>\alpha$ and we have a contradiction because $\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\mid H_{S}]\leq\alpha$.
If $\alpha^{\mathrm{Perm}}=\alpha$, then $\phi^{\mathrm{Simes}}_{S}=\phi^{\mathrm{PermSimes}}_{S}$. By Corollary 3.5 in Marandon et al. (2024), if $\alpha(m+1)/|S|$ is an integer, then all inequalities in (1) becomes equalities. ∎
Table A2 below presents the size of the tests $\phi^{\mathrm{Simes}}$ and $\phi^{\mathrm{SimesPerm}}$, along with the critical value $\alpha^{\mathrm{Perm}}$, as a function of $m$, with $\alpha=0.1$ and $n=3$.
$m$ 9 14 19 24 29 34 39 44 49 54
$\mathbb{P}(\phi_{S}^{\mathrm{Simes}}=1|H_{S})$ 0.000 0.022 0.013 0.009 0.100 0.086 0.075 0.072 0.064 0.058
$\mathbb{P}(\phi_{S}^{\mathrm{SimesPerm}}=1|H_{S})$ 0.055 0.025 0.014 0.009 0.100 0.097 0.083 0.072 0.065 0.058
$\alpha^{\mathrm{Perm}}$ 0.200 0.133 0.100 0.080 0.100 0.171 0.150 0.111 0.100 0.091
Table A2: Size of the tests $\phi^{\mathrm{Simes}}$ and $\phi^{\mathrm{SimesPerm}}$, along with the critical value $\alpha^{\mathrm{Perm}}$, as a function of $m$, with $\alpha=0.1$ and $n=3$.
A1.3 The Fisher Combination Test for Conformal p-Values
An alternative classical approach for testing $H_{S}$ based on conformal p-values $(p_{j})_{j\in S}$ is provided by the Fisher combination method (Fisher, 1925). This method originally assumed independence but was later refined by Bates et al. (2023) to accommodate the (weak) positive dependencies exhibited by conformal p-values. The idea is to evaluate the statistic $T^{\mathrm{Fisher}}_{S}=-2\sum_{j\in S}\log(p_{j})$ and then reject $H_{S}$ at level $\alpha$ if and only if $T^{\mathrm{Fisher}}_{S}>c^{\chi^{2}}_{\alpha}(|S|)\sqrt{1+|S|/m}-2|S|(\sqrt{1+|S|/m}-1)$, where $c^{\chi^{2}}_{\alpha}(|S|)$ is the $(1-\alpha)$-quantile of the $\chi^{2}$ distribution with $2|S|$ degrees of freedom. The indicator of this rejection event can thus be written as:
$\phi^{\mathrm{Fisher}}_{S}$ $=\mathds{1}\left\{-2\sum_{j\in S}\log(p_{j})>c^{\chi^{2}}_{\alpha}(|S|)\sqrt{1+|S|/m}-2|S|\left(\sqrt{1+|S|/m}-1\right)\right\}.$ (A15)
Bates et al. (2023) proved that this test is asymptotically valid in the limit of large $|S|$ and $m$, with both sample sizes growing at the same rate. This differs from the classical Fisher test (Fisher, 1925), which assumes the $p$-values are mutually independent and can be recovered from (A15) by letting $m\to\infty$ while holding $|S|$ fixed.
A1.4 The Wilcoxon-Mann-Whitney Rank Test
An important class of permutation tests is defined in terms of sample ranks, known as rank tests. These tests rely on the ranks of the data, making them invariant to monotone transformations, as these transformations don’t alter the relative ranking of observations.
For any $S\subseteq[n]$ and any $j\in S$, let $R_{j}$ denote the rank of $Y_{j}$ among $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. The dependence of $R_{j}$ on $S$ is left implicit to simplify the notation, since this does not create ambiguity.
This classical two-sample Wilcoxon (Wilcoxon, 1945) or Mann-Whitney (Mann and Whitney, 1947) test rejects $H_{S}$ for large values of the sum of ranks in the test sub-sample $S$
$T^{\mathrm{W}}_{S}$ $=\sum_{j\in S}R_{j},$ (A16)
or for large values of the U-statistic
$T^{\mathrm{MW}}_{S}$ $=\sum_{j\in S}\sum_{i=1}^{m}\mathds{1}\{X_{i}<Y_{j}\}.$ (A17)
The Mann Whitney U-statistic can be expressed as $T^{\mathrm{MW}}_{S}=T^{\mathrm{W}}_{S}+|S|(|S|+1)/2$, differing only by a constant term. The two formulations result in the same test, which we will call the Wilcoxon-Mann-Whitney (WMW) test. The WMW level-$\alpha$ test is given by
$\phi^{\mathrm{WMW}}_{S}$ $=\mathds{1}\left\{T^{\mathrm{W}}_{S}\geq c^{\text{W}}_{\alpha}(m,|S|)\right\}=\mathds{1}\left\{T^{\mathrm{MW}}_{S}\geq c^{\text{MW}}_{\alpha}(m,|S|)\right\}.$ (A18)
The critical values $c^{\text{W}}_{\alpha}(m,|S|)$ and $c^{\text{MW}}_{\alpha}(m,|S|)$ are the $(1-\alpha)$ quantiles of the permutation distribution of $T^{\mathrm{W}}_{S}$ and $T^{\mathrm{MW}}_{S}$, respectively. Since the two test statistics result in the same test, we will denote them generally by $T^{\mathrm{MWW}}_{S}$ and the critical value by $c^{\text{WMW}}_{\alpha}(m,|S|)$.
For small samples, the permutation null distribution of $T^{\text{WMW}}_{S}$ can be found either via recursion (Mann and Whitney, 1947) or direct permutations. The test $\phi^{\mathrm{WMW}}_{S}$ is of exact size $\alpha$, i.e. $\mathbb{P}(\phi^{\mathrm{WMW}}_{S}=1|H_{S})=\alpha$, for $\alpha\in\Lambda=\{a_{r},r\in[m|S|]\}$, where $a_{r}=\mathbb{P}(T^{\mathrm{WMW}}_{S}\geq r|H_{S})$; for $\alpha\in(0,1)\setminus\Lambda$, it is conservative, i.e. $\mathbb{P}(\phi^{\mathrm{WMW}}_{S}=1|H_{S})<\alpha$.
A well-known approach when $m$ and $|S|$ are both large is that of Hoeffding (1948), which is based on an application of the Central Limit Theorem for U-statistics. Under the null hypothesis $H_{S}$, for large $m$ and $|S|$, the Mann Whitney statistic $T^{\text{MW}}_{S}$ is approximately normally distributed:
$T^{\text{MW}}_{S}$ $\approx N\left(\frac{|S|m}{2},\frac{m|S|(m+|S|+1)}{12}\right).$ (A19)
Other asymptotic approximations include the Edgeworth expansion of Fix and Hodges (1955) and the uniform approximation of Buckle et al. (1969).
It is also interesting to note that the Mann-Whitney statistic $T^{\text{MW}}_{S}$ reduces to $T^{\text{MW}}_{j}=\sum_{i=1}^{m}\mathds{1}\{X_{i}<Y_{j}\}$ if $S=\{j\}$, for any $j\in[n]$. The statistic $T^{\text{MW}}_{j}$ can be equivalently expressed as a function of the conformal p-value $p_{j}$ in  (3):
$T^{\mathrm{MW}}_{j}$ $=\sum_{i=1}^{m}\mathds{1}\{X_{i}<Y_{j}\}=m-\sum_{i=1}^{m}\mathds{1}\{X_{i}\geq Y_{j}\}=(m+1)(1-p_{j})$ (A20)
When $S=\{j\}$, the WMW two-sample test, with the 1st sample being $X_{1},\ldots,X_{n}$ and the 2nd “sample” being just $Y_{j}$, rejects $H_{j}$ if the rank of $Y_{j}$ in the sequence $(X_{1},\ldots,X_{n},Y_{j})$ is large or, equivalently, if the conformal p-value $p_{j}$ in  (3) is small. Therefore, the WMW test provides a particularly intuitive bridge between the modern framework of conformal inference and the classical world of two-sample rank tests (Kuchibhotla, 2020).
Another interesting connection, formally stated in the next proposition, is that the Mann-Whitney test for $H_{S}$ can be defined by the average of conformal p-values test statistic, also known as Edgington’s method (Edgington, 1972).
Proposition A2.
The Mann-Whitney test statistic $T^{\text{MW}}_{S}$ can be equivalently expressed as a function of the conformal p-values $p_{j}$ in  (3):
$T^{\text{MW}}_{S}$ $=(m+1)\sum_{j\in S}(1-p_{j}).$
Then the WMW test can be defined by the average of conformal p-values test statistic, also known as Edgington’s method (Edgington, 1972):
$\bar{p}_{S}$ $=\frac{1}{|S|}\sum_{j\in S}p_{j}.$
Proof.
Note that The Mann-Whitney test statistic $T^{\text{MW}}_{S}$ is the sum of the individual contributions $T^{\text{MW}}_{j}$ with $j\in S$, i.e. $T^{\text{MW}}_{S}=\sum_{j\in S}T^{\text{MW}}_{j}$. Then $T^{\text{MW}}_{S}=(m+1)\sum_{j\in S}(1-p_{j})$ follows from  (A20). Since $T^{\text{MW}}_{S}$ is a monotonic function of $\bar{p}_{S}=|S|^{-1}\sum_{j\in S}p_{j}$, the test statistic can be simplified to the average of conformal p-values $\bar{p}_{S}$, leading to rejection of $H_{S}$ for small values of $\bar{p}_{S}$. ∎
Appendix A2 Estimating the Outlier Distribution
This section details how to empirically estimate the outlier distribution $G$, which is essential for implementing a data-driven version of Shirashi’s LMPI local testing procedure as described in Section 2.3.2. Recall from Section 2.3.2 that Shirashi’s optimality result (Shiraishi, 1985) is based on the following mixture model:
[…]
where $\theta\in[0,1]$ represents the proportion of outliers, and $G$ is a distribution function on $[0,1]$. In the following, we describe a practical and effective approach for computing an estimate $\hat{G}$ of $G$ using the data in $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$, while ensuring that an approximately valid and approximately LMPI test can still be performed using Shirashi’s method conditional on $\hat{G}$.
A2.1 Diluting the Test Set for Conditional Exchangeability
The key idea, inspired by approaches proposed in different contexts by Liang et al. (2024) and Marandon et al. (2024), is to randomly split the reference sample $(X_{1},\ldots,X_{m})$ into two disjoint subsets, namely $(X_{1},\ldots,X_{m_{1}})$ and $(X_{m_{1}+1},\ldots,X_{m})$, for some choice of $m_{1}$ such that $1<m_{1}<m$, like $m_{1}=\lceil m/2\rceil$. Intuitively, $(X_{1},\ldots,X_{m_{1}})$ is used as a smaller reference set and $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$ as a “diluted” test set when fitting $G$. More precisely, the estimation of $G$ is carried out by fitting the following mixture model:
[…]
where $\theta_{1}=(\theta n+m_{1})/(n+m_{1})$ and $G$ is the same outlier distribution function as in Equation (A21). Any approach may be applied to fit this mixture model, as long as it is invariant to the ordering of the data in $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$. A concrete approach is discussed in the next subsection.
The motivation for this method is that it ensures the inlier scores among $(Y_{1},\ldots,Y_{n})$ remain exchangeable with the smaller reference set $(X_{m_{1}+1},\ldots,X_{m})$ conditional on the estimated $\hat{G}$. Therefore, any downstream conformal inferences that utilizes $(X_{m_{1}+1},\ldots,X_{m})$ as a reference (or calibration) set are still valid. This is akin to the idea underlying AdaDetect, the individual outlier identification method proposed by Marandon et al. (2024).
A subtle point worth highlighting is that the asymptotic critical values for Shirashi’s LMPI test outlined in Section 2.3.2 were derived under the assumption that all inliers are i.i.d. (A21), whereas conditional on $\hat{G}$ the inliers among $(Y_{1},\ldots,Y_{n})$ are only exchangeable with $(X_{m_{1}+1},\ldots,X_{m})$ but not independent. However, to the best of our knowledge, this issue does not seem to have noticeable practical implications, even for only moderately large sample sizes. We refer to Section 2.6 for a more detailed discussion of this point.
A2.2 An Intuitive Mixture Modeling Approach
Although there is an extensive literature on fitting mixture models like the one in (A22), covering a wide range of parametric and nonparametric methods, a full review is beyond our scope. Instead, we present a simple and intuitive approach that has proven effective for our purposes.
  1. 1.
    Rescale the data, $X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n}$, to be between 0 and 1, if they are not already scaled as such (many standard classifiers output scores in this range by default).
  2. 2.
    Fit a $\text{Beta}(b_{1},b_{2})$ distribution to $X_{1},\ldots,X_{m_{1}}$ via maximum likelihood. To avoid numerical instabilities from values too close to 0 or 1, threshold all data points between 0.001 and 0.999. Let $\hat{b}_{1}$ and $\hat{b}_{2}$ be the estimated parameters.
  3. 3.
    Apply an inverse CDF transform based on the fitted $\text{Beta}(\hat{b}_{1},\hat{b}_{2})$ distribution to $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$, making the inlier distribution approximately uniform on $[0,1]$. Let $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ denote the transformed test set.
  4. 4.
    Fit the following mixture model via maximum likelihood:
[…]
    where $G$ is approximated by a $\text{Beta}(b^{\prime}_{1},b^{\prime}_{2})$ distribution. The estimated parameters are denoted as $\hat{b}^{\prime}_{1}$ and $\hat{b}^{\prime}_{2}$.
  5. 5.
    Output the estimated outlier distribution: $\hat{G}\approx\text{Beta}(\hat{b}^{\prime}_{1},\hat{b}^{\prime}_{2})$.
While this approach may seem somewhat heuristic, as it is not aimed at making rigorous inferences about $G$, it is sufficient for our purposes. The invariance of this estimation procedure to the ordering of $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ ensures the validity of our test regardless of how closely $\hat{G}$ approximates the true $G$. Moreover, as demonstrated in our numerical experiments, this estimation method performs well enough in practice to provide our testing method with a noticeable advantage over standard approaches, such as the WMW test, particularly in scenarios where the oracle LMPI test has a significant edge.
A2.3 Enabling Closed Testing Shortcuts for Monotone Statistics
When applied within a closed testing framework, Shirashi’s LMPI testing procedure can become computationally expensive without an appropriate shortcut. As discussed in Appendix A3, a fast and exact shortcut exists, but it relies on the monotonicity of the outlier probability density—specifically, the derivative of the function $\hat{G}$ estimated above. This motivates the need for an estimation procedure that guarantees a monotone outlier density.
Although more sophisticated approaches have been studied to address this issue (Patra and Sen, 2016), a simple post-hoc solution that works well for our purposes is to approximate the probability density function of the $\text{Beta}(\hat{b}^{\prime}_{1},\hat{b}^{\prime}_{2})$ distribution, obtained as described in the previous section, with a monotone increasing function using isotonic regression on a finite grid of values. Since it may not be clear in advance whether the best approximation of the outlier density is increasing or decreasing, the direction of monotonicity is adaptively chosen by comparing the residual sum of squares for the two corresponding isotonic regression models.
While this method is somewhat heuristic, we have found it to be more reliable in practice than other, more sophisticated approaches we tried. Nonetheless, it is likely that this approach could be further refined with some additional effort.
Appendix A3 Computational Shortcuts for Closed Testing
In general, computing $d(S)$ in (8) involves the evaluation of exponentially many tests, which hinders its practical application. For specific local tests, however, there exist polynomial-time shortcuts (Goeman et al., 2019, 2021; Tian et al., 2023).
A3.1 Shortcuts for Simes Local Tests
With Simes’ local test, the shortcut described in Goeman et al. (2019) and given in Algorithm 4 allows calculating $d_{\mathrm{Simes}}(S)$ in linear time, after an initial preparatory step that takes $\mathcal{O}(n\log n)$ time.
Input : $p$-values $p_{1},\ldots,p_{n}$; index set $S\subseteq[n]$; significance level $\alpha\in(0,1)$.
1
2Sort the $p$-values $p_{(1)}\leq\ldots\leq p_{(n)}$;
3Compute $h=\max\{0\leq k\leq n:p_{(n-k+j)}>j\alpha/k\,\,\mathrm{for\,\,}j=1,\ldots,k\}$;
4Compute $d_{\mathrm{Simes}}(S)=\min\{0\leq k\leq|S|:p_{(k+j:S)}>j\alpha/h\,\,\mathrm{for\,\,}j=1,\ldots,|S|-k\}$;
Output : A simultaneous $(1-\alpha)$-confidence lower bound $d_{\mathrm{Simes}}(S)$ for the number of outliers in $S$, for any $S\subseteq[n]$.
5
Algorithm 4 Shortcut for closed testing with Simes’ test (Goeman et al., 2019)
Line 2 of Algorithm 4 gives $h$, which is a $(1-\alpha)$-confidence upper bound for the number of true inliers $|I_{0}|$ in the overall test set. Then $d_{\mathrm{Simes}}=n-h$ is a $(1-\alpha)$-confidence lower bound for the number of false hypotheses (cfr. number of true outliers in the test sample) $|I_{1}|$.
The index set of the discoveries (cfr. localized outliers) is given by $D_{\mathrm{Simes}}=\{j\in[n]:d_{{\mathrm{Simes}}}(\{j\})=1\}=\{j\in[n]:p_{j}\leq\alpha/h\}.$ All hypotheses $H_{j}$ with index $j\in D_{\mathrm{Simes}}$ are rejected by closed testing while controlling the familywise error rate (FWER) at level-$\alpha$ (Hommel, 1988): $\mathbb{P}(|D_{\mathrm{Simes}}\cap I_{0}|>0)\leq\alpha$.
A3.2 Closed Testing with Simes Local Tests vs BH
The Simes’ test and the Benjamini-Hochberg (BH) procedure share the same critical values, and we can expect some relationships between the two methods.
Let $\phi$ denote a level-$\alpha$ test for the global null hypothesis $H=H_{[n]}$. Closed testing provides the following quantities:
  • $\mathds{1}\{\phi=1\}$ to indicate outlier detection;
  • $d$ to provide outlier enumeration;
  • $|D|$ to quantify outlier discovery.
Outlier enumeration involves counting the number of true outliers in the test sample without necessarily pinpointing their exact locations. Outlier enumeration is implied by outlier discovery (by using $|D|$) and, in turn, implies outlier detection (by using $\mathds{1}\{d>0\}$). These quantities can be applied to any subset $S$ of the test sample, obtaining $\mathds{1}\{\phi_{S}=1\}$, $d(S)$, and $|D\cap S|$. Simultaneous confidence bounds $d(S)$ obtained by closed testing satisfy the following relations:
$\mathds{1}\{\phi_{S}=1\}\geq\mathds{1}\{d(S)>0\},\qquad d(S)\geq|D\cap S|.$ (A24)
The BH algorithm applied to $p_{1},\ldots,p_{n}$ at level $\alpha$ returns the index set of the discoveries:
$D_{\mathrm{BH}}$ $=$ $\{j\in[n]:p_{j}\leq\alpha d_{\mathrm{BH}}/n\}$ (A25)
where
$d_{\mathrm{BH}}$ $=$ $\max\left\{k\in\{0,\ldots,n\}:\sum_{j=1}^{n}\mathds{1}\{p_{j}\leq\alpha k/n\}\geq k\right\}.$ (A26)
is the number of discoveries made by the BH procedure. To distinguish between discoveries with FWER and FDR guarantees, we will refer to discoveries made by closed testing as FWER discoveries, while those made by an FDR controlling procedure will be called FDR discoveries.
The Simes’ test for $H$ and the BH procedure share the same critical values, and we can expect some relationships between the two methods. A simple observation is that the event where the CTSimes and the BH procedure make at least one non-trivial statement is identical:
$\mathds{1}\{d_{\mathrm{Simes}}>0\}=\mathds{1}\{d_{\mathrm{BH}}>0\}.$ (A27)
Consequently, the two methods have the same weak FWER control and the same power for outlier detection. Furthermore, Goeman et al. (2019) showed that the confidence bound $d_{\mathrm{Simes}}$ is between the number of Hommel-FWER discoveries and the number of BH-FDR discoveries, i.e.
$|D_{\mathrm{Simes}}|\leq d_{\mathrm{Simes}}\leq|D_{\mathrm{BH}}|.$ (A28)
Finally, note that the BH discoveries do not provide the true discovery guarantee, i.e.,
$\mathbb{P}(|D_{\mathrm{BH}}\cap S|\leq|I_{1}\cap S|\mathrm{\,\,for\,\,all\,\,}S)$ (A29)
may be less than $1-\alpha$.
A3.3 Shortcuts for Storey-Simes Local Tests
The BH procedure guarantees that expected proportion of inliers among the discoveries is bounded by $\alpha\pi_{0}$, where $\pi_{0}=|I_{0}|/n$ is the unknown proportion of inliers in the test sample. When $\pi_{0}$ is expected to be not close to 1, $\alpha\pi_{0}$ falls below the target level $\alpha$, making the BH procedure slightly too conservative. This motivates adjusting the level into $\alpha/\hat{\pi}_{0}$ with $\hat{\pi}_{0}$ an estimate of $\pi_{0}$, resulting in a $\pi_{0}$-adaptive version of the BH algorithm.
The idea of the adaptive BH procedure extends to the Simes test, leading to a $\pi_{0}$-adaptive version of Simes’ test for $H_{S}$ that incorporates the proportion of true null hypotheses in $S$. The test is formulated as:
$\phi_{S}^{\mathrm{Storey-Simes}}=\mathds{1}\left\{\min_{k\in\{1,\ldots,|S|\}}\{|S|p_{(k:S)}/k\}\leq\alpha/\hat{\pi}_{0}^{S}\right\},$ (A30)
where
$\hat{\pi}^{S}_{0}$ $=$ $\frac{1+\sum_{j\in S}\mathds{1}\{p_{j}>\lambda\}}{|I|(1-\lambda)}$ (A31)
and $\lambda=h/(|S|+1)$ for any pre-specified integer $h$. This estimator is known as the Schweder-Spjøtvoll or Storey’s estimator (Schweder and Spjøtvoll, 1982; Storey, 2002; Storey et al., 2004). The test $\phi_{I}^{\mathrm{Storey-Simes}}$ was considered in Bogomolov (2023) and Heller and Solari (2023).
The shortcut given in Algorithm 5 allows calculation of $d_{\mathrm{Storey-Simes}}(S)$ in quadratic time.
Input : $p$-values $p_{1},\ldots,p_{n}$; index set $S\subseteq[n]$; significance level $\alpha\in(0,1)$; tuning parameter $\lambda\in(0,1)$.
1
2Initialize $d_{\mathrm{Storey-Simes}}(S)\leftarrow 0$;
3for each $i\in\{1,\ldots,|S|\}$ do
4       for each $j\in\{0,\ldots,n-|S|+i-1\}$ do
5
6            $l\leftarrow|S|-i+j+1$;
7            $(q_{1},\ldots,q_{l})\leftarrow(p_{(i:S)},\ldots,p_{(|S|:S)},p_{(n-|S|-j+1:S^{c})},\ldots,p_{(n-|S|:S^{c})})$;
8            if  $\min_{j\in[l]}\Big{(}l\frac{q_{(j)}}{j}\Big{)}>\alpha\Big{(}\frac{l(1-\lambda)}{1+\sum_{j\in[l]}\mathds{1}\{q_{j}>\lambda\}}\Big{)}$,  then
9                  return $d_{\mathrm{Storey-Simes}}(S)$.
10
11
12      $d_{\mathrm{Storey-Simes}}(S)=d_{\mathrm{Storey-Simes}}(S)+1$;
13
Output : $(1-\alpha)$-confidence lower bound $d_{\mathrm{Storey-Simes}}(S)$ for the number of true discoveries in $S$
Algorithm 5 Shortcut for closed testing with adaptive Simes local tests.
The closed testing with adaptive Simes’ test ensures the true discovery guarantee by observing that the adaptive Simes’ test rejects $H_{I_{0}}$ if and only if the adaptive Benjamini-Hochberg procedure applied to conformal $p$-values $(p_{j},j\in I_{0})$ rejects at least one hypothesis. This rejection occurs with a probability bounded by $\alpha$, as shown in Corollary 2.5 and Theorem 2.6 in Bates et al. (2023), and Theorem 3.1 in Mary and Roquain (2022).
A3.4 Shortcut for Shiraishi Local Tests
The shortcut provided in Algorithm 6 enables the calculation of $d_{\mathrm{Shiraishi}}(S)$ in quadratic time for a given increasing density function $g$. Let $a^{l}=(a^{l}_{1},\ldots,a^{l}_{m+l})$ be a vector where the $k$th element is $a^{l}_{k}=\mathbb{E}[g(U^{(k)}_{m+l})]$. Since $g$ is increasing, the elements of the vector are ordered such that $a_{1}^{l}\leq\ldots\leq a^{l}_{m+l}$. The score vector $a^{l}$ can be estimated with the desired accuracy using Monte Carlo simulation. The critical value $c^{G}_{\alpha}(m,l)$ corresponds to the $(1-\alpha)$-quantile of the normal distribution with the following mean and variance:
$\mu_{m+l}=\frac{l}{m+l}\sum_{k\in[m+l]}a^{l}_{k},\qquad\sigma^{2}_{m+l}=\frac{ml}{(m+l)(m+l-1)}\sum_{k\in[m+l]}\left(a^{l}_{k}-\frac{\mu_{m+l}}{l}\right)^{2}.$
Input : Calibration scores $(X_{1},\ldots,X_{m})$; test scores $(Y_{1},\ldots,Y_{n})$; an increasing density function $g$; index set $S\subseteq[n]$; significance level $\alpha\in(0,1)$.
1
2Initialize $d_{\mathrm{Shiraishi}}(S)\leftarrow 0$;
3for each $i\in\{1,\ldots,|S|\}$ do
4       for each $j\in\{0,\ldots,n-|S|+i-1\}$ do
5
6            $l\leftarrow|S|-i+j+1$;
7            $(W_{1},\ldots,W_{l})\leftarrow(Y_{(1:S)},\ldots,Y_{(|S|-i+1:S)},Y_{(1:S^{c})},\ldots,Y_{(j:S^{c})})$;
8            Let $R_{j}$ denote the rank of $W_{j}$ among $(X_{1},\ldots,X_{m},W_{1},\ldots,W_{l})$;
9            if  $\sum_{j=1}^{l}\mathbb{E}[g(U_{m+l}^{(R_{j})})]<c^{G}_{\alpha}(m,l)$,  then
10                  return $d_{\mathrm{Shiraishi}}(S)$.
11
12
13      $d_{\mathrm{Shiraishi}}(S)=d_{\mathrm{Shiraishi}}(S)+1$;
14
15
Output : A $(1-\alpha)$-lower bound $d_{\mathrm{Shiraishi}}(S)$ for the number of outliers in $S$.
16
Algorithm 6 Shortcut for closed testing with Shiraishi local tests.
A3.5 Shortcut for Monotone, Symmetric, and Separable Tests
Tian et al. (2023) introduced a shortcut for the general case of closed testing with local tests satisfying three properties: monotonicity, symmetry and separability. These properties are met by classic Mann-Whitney test and Fisher’s combination test. For precise definitions of these properties, please refer to Appendix B in Tian et al. (2023).
Algorithm 1 in Tian et al. (2023) enables a linear-time computation of the $(1-\alpha)$-confidence lower bound $d(S)$ for any $S\subseteq[n]$, after a preparatory step involving the computation and sorting of the test statistics.
Appendix A4 Additional Empirical Results
A4.1 Numerical Experiments with Synthetic Data
Figures A1 and A2 describe findings similar to those in Figures 2 and 3, respectively. The distinction is that now the data are simulated from a binomial model borrowed from Liang et al. (2024), for which binary classifiers are more powerful than one-class classifiers. Figures A3 and A4 summarize the 90-th empirical quantiles of the lower confidence bounds for the numbers of outliers presented in Figures 2 and A1, respectively. Figure A5 delves into experiments related to those depicted in Figures A1 and A2, but utilizing synthetic data from a mixture model that bridges between the distributions considered above. These results highlight ACODE’s flexibility, which selects an effective classifier and local testing procedure for each specific case.
Figures A6 and A7 offer additional insights into the performance of the LMPI local testing procedure, outlined in Section 2.3.2, when applied in different settings. These results highlight its practical advantages in situations where the outlier scores are either underdispersed or overdispersed relative to the inlier scores.
Figure A1: Median values for a 90% lower confidence bound on the total number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. The results are shown as a function of the true number of outliers within a test set of size 1000. In these experiments, the synthetic data are generated from a binomial model borrowed from Liang et al. (2024).
Figure A2: Lower confidence bounds for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
Figure A3: Empirical 90-th quantile for a 90% lower confidence bounds on the total number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. Other details are as in Figure 2.
Figure A4: Empirical 90-th quantile for a 90% lower confidence bounds on the total number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. Other details are as in Figure A1.
Figure A5: Lower confidence bounds for the total number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. The results are shown as a function of the true number of outliers within a test set of size 1000. In these experiments, the synthetic data are generated from a binomial model borrowed from Liang et al. (2024). Other details are as in Figure A1.
Figure A6: Power of different rank tests for the global null hypothesis of no outliers in a test set containing 200 independent observations of a simulated univariate score, using a reference set containing 500 independent inlier scores. The nominal level is $\alpha=0.05$ (horizontal dashed line). The results are shown as a function of the true number of outliers in the test set. Top: the inliers are randomly sampled from a uniform distribution on $[0,1]$, while the outliers are sampled from the alternative distribution indicated in each panel. Bottom: the inliers are standard normal, while the outlier scores follow a non-standard normal distribution, as indicated in each panel. Three versions of the LMPI test from Section 2.3.2 are compared. The first version uses oracle knowledge of the true transformation $G$ linking the outlier distribution to the inlier distribution. The second version of the LMPI test uses an empirical estimate of $G$ obtained as described in Appendix A2. The third version relies on a monotone approximation of the derivative $g=G^{\prime}$, which (in Figure A7) enables a computationally convenient shortcut in the context of closed testing, as also explained in Appendix A2.
Figure A7: Median values for a 90% lower confidence bound on the total number of outliers in the test set, in the same experiments of Figure A6. The lower bound is calculated through closed testing, using different local testing procedures. For the LMPI approach, if the derivative $g$ of the oracle function $G$ is not monotone, it is in practice replaced by a monotone approximation that enables the application of a computationally efficient closed testing shortcut.
A4.2 Numerical Experiments with Particle Collision Data
Figure A8 is a more detailed version of Figure 1, including a broader range for the choice of the local test. Figures A9 and A10 report on experiments similar to those of Figures 1 and 5, respectively, but leveraging a larger training set containing 100,000 inliers, which leads to higher power for all methods. Tables A3 and A4 provide a more detailed view of the results, respectively, in Figures A8 and A9. Finally, Figures A11 and A12 summarize the 90-th empirical quantiles of the lower confidence bounds for the numbers of outliers presented in Figures 5 and A10, respectively, confirming again the validity of our method.
Figure A8: Performance of ACODE for collective outlier detection with the LHCO data, in the experiments of Figure 1. Our method utilizes a local testing procedure that may be adaptively selected (red curve) or fixed (other solid curves). Compared to Figure 1, these results include a more detailed comparison of the performance of ACODE applied using different choices of local testing procedures.
Local testing procedure
Outliers Simes Storey-Simes Fisher WMW LMPI (Lehmann, k=3) Adaptive Cherry picking
90% Lower bound (median)
   0 0 (0) 0 (0) 0 (2) 0 (5) 0 (4) 0 (5) 0 (4)
   200 0 (0) 0 (0) 0 (1) 0 (5) 0 (4) 0 (5) 0 (5)
   500 0 (0) 0 (0) 94 (11) 82 (17) 123 (16) 147 (17) 224 (14)
   1000 4 (12) 39 (13) 593 (24) 640 (29) 665 (26) 684 (28) 736 (14)
   1500 281 (20) 337 (19) 1055 (23) 1106 (27) 1137 (25) 1164 (24) 1190 (14)
90% Lower bound (90-th quantile)
   0 0 (0) 0 (0) 0 (2) 0 (5) 0 (4) 0 (5) 27 (4)
   200 0 (0) 0 (0) 14 (1) 28 (5) 42 (4) 82 (5) 104 (5)
   500 0 (0) 0 (0) 227 (11) 310 (17) 301 (16) 341 (17) 404 (14)
   1000 188 (12) 198 (13) 700 (24) 871 (29) 840 (26) 864 (28) 910 (14)
   1500 458 (20) 501 (19) 1188 (23) 1341 (27) 1347 (25) 1358 (24) 1376 (14)
Power (global null)
   0 0.00 (0.00) 0.00 (0.00) 0.05 (0.03) 0.09 (0.04) 0.09 (0.04) 0.09 (0.04) 0.15 (0.04)
   200 0.00 (0.00) 0.00 (0.00) 0.17 (0.05) 0.14 (0.04) 0.17 (0.05) 0.18 (0.05) 0.33 (0.05)
   500 0.01 (0.01) 0.01 (0.01) 0.72 (0.06) 0.69 (0.06) 0.78 (0.05) 0.79 (0.05) 0.88 (0.03)
   1000 0.52 (0.07) 0.55 (0.06) 0.97 (0.02) 0.98 (0.02) 1.00 (0.00) 0.98 (0.02) 1.00 (0.00)
   1500 0.95 (0.03) 0.97 (0.02) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Table A3: Performance of ACODE for collective outlier detection with the LHCO data, as a function of the true number of outliers. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). The numbers in parenthesis are standard errors. Other details are as in Figures 1 and A8.
Figure A9: Performance of ACODE for collective outlier detection with the LHCO data, using a training set containing 100,000 inliers. In the interest of computational efficiency, in these experiments, the conformity scores leveraged by ACODE are computed by a fixed AdaBoost classifier. Other details are as in Figure 1.
Local testing procedure
Outliers Simes Storey-Simes Fisher WMW LMPI (Lehmann, k=3) Adaptive Cherry picking
90% Lower bound (median)
   0 0 (0) 0 (0) 0 (1) 0 (3) 0 (2) 0 (2) 0 (3)
   200 0 (0) 0 (0) 0 (4) 0 (7) 0 (7) 0 (7) 0 (8)
   500 0 (0) 0 (0) 204 (11) 202 (16) 247 (15) 273 (15) 284 (15)
   1000 88 (10) 118 (10) 666 (10) 720 (18) 751 (14) 758 (13) 796 (13)
   1500 343 (16) 392 (15) 1114 (10) 1224 (20) 1242 (16) 1252 (15) 1276 (15)
90% Lower bound (90-th quantile)
   0 0 (0) 0 (0) 0 (1) 0 (3) 0 (2) 2 (2) 8 (3)
   200 0 (0) 0 (0) 46 (4) 107 (7) 131 (7) 133 (7) 156 (8)
   500 0 (0) 0 (0) 328 (11) 421 (16) 436 (15) 456 (15) 474 (15)
   1000 238 (10) 251 (10) 782 (10) 952 (18) 919 (14) 932 (13) 957 (13)
   1500 541 (16) 563 (15) 1255 (10) 1467 (20) 1430 (16) 1452 (15) 1478 (15)
Power (global null)
   0 0.00 (0.00) 0.00 (0.00) 0.04 (0.02) 0.06 (0.02) 0.07 (0.03) 0.11 (0.03) 0.14 (0.03)
   200 0.00 (0.00) 0.00 (0.00) 0.28 (0.05) 0.23 (0.04) 0.31 (0.05) 0.35 (0.05) 0.40 (0.05)
   500 0.06 (0.02) 0.07 (0.03) 0.89 (0.03) 0.78 (0.04) 0.84 (0.04) 0.88 (0.03) 0.89 (0.03)
   1000 0.76 (0.04) 0.83 (0.04) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
   1500 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Table A4: Performance of ACODE for collective outlier detection with the LHCO data, using a training set containing 100,000 inliers. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis (bottom). The numbers in parenthesis are standard errors. Other details are as in Figure A9.
Figure A10: Lower confidence bounds for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve corresponds to the true number of outliers in this selected set. In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
Figure A11: Empirical 90-th quantile for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in the same numerical experiments of Figure 5.
Figure A12: Empirical 90-th quantile for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in the same numerical experiments of Figure A10.
A4.3 Additional Experiments with Real Data
In this section, we further investigate the performance of ACODE by applying it to 6 additional data sets previously utilized in the related literature. These data sets are:
  • CreditCard (Dataset, c), containing 30 variables, 492 outliers, 284315 inliers;
  • Pendigits (Dataset, e), containing 16 variables, 156 outliers, 6714 inliers;
  • Covertype (Dataset, b), containing 10 variables, 2747 outliers, 286048 inliers;
  • Shuttle (Dataset, f), containing 9 variables, 3511 outliers, 45586 inliers;
  • Mammography (Dataset, d), containing 6 variables, 260 outliers, 10923 inliers;
  • ALOI (Dataset, a), containing 27 variables, 1508 outliers, 48026 inliers.
These experiments are conducted following an approach similar to that described in Section 3.1, applying our method based on the same suite of 6 classification algorithms and 6 local testing procedures. Each experiment is independently repeated 100 times, each time randomly sampling disjoint training, calibration, tuning, and test sets of sizes 1000, 100, 100, and 100, respectively. The training, calibration, and tuning sets include only inliers, while the proportion of outliers in the test set is varied as a control parameter between 0 and 1.
Figures A13A18 summarize, separately for each data set, the performances of the lower confidence bounds and global tests obtained with our method. The results are reported as a function of the true number of outliers in the test set and are also stratified based on the type of classifier utilized by ACODE and on the underlying local testing procedure. The findings overall indicate that binary classification models generally yield higher power in these experiments, which can be attributed to the relatively low dimensionality of these data sets. Moreover, among the local testing methods considered, both Fisher’s method and the WMW method demonstrate similarly high power, typically outperforming Simes’ method. Crucially, the fully automatic implementation of ACODE, which selects both the classification model and the local testing procedure in a data-adaptive way, leads to near-oracle performance in all cases.
Finally, Figures A19 and A20 describe experiments aimed at constructing 90% lower confidence bounds for the number of outliers within a data-dependent subset of test points, specifically those selected for their high conformity scores. For easier understanding of these experiments, ACODE is now applied with a fixed classification algorithm, the one-class isolation forest, rather than leveraging the full suite of 6 different algorithms. The results show that the WMW test consistently outperforms the Fisher combination method in terms of power, aligning with findings previously documented in Sections 3.1 and 3.2. Furthermore, the fully adaptive implementation of ACODE effectively maximizes power by automatically identifying the most powerful local testing procedure for each case, again achieving oracle-like performance.
Figure A13: Performance of ACODE for collective outlier detection with the creditcard data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure 2.
Figure A14: Performance of ACODE for collective outlier detection with the pendigits data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure A13.
Figure A15: Performance of ACODE for collective outlier detection with the covertype data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure A13.
Figure A16: Performance of ACODE for collective outlier detection with the shuttle data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure A13.
Figure A17: Performance of ACODE for collective outlier detection with the mammography data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure A13.
Figure A18: Performance of ACODE for collective outlier detection with the aloi data set, as a function of the true number of outliers in a test set of size 100. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). Other details are as in Figure A13.
Figure A19: Median values for a lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments with several different data sets. The results are shown as a function of the proportion of selected test points. In these experiments, ACODE is applied using a one-class isolation forest model to compute the conformity scores. Other details are as in Figure 3.
Figure A20: Empirical 90-th quantile for a lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments with several different data sets. Other details are as in Figure A19.